Wednesday, October 18, 2017

From HallCWiki
Jump to: navigation, search

Meeting Information

  • Location CC L210A, 11 am
  • DocDB

Blue Jeans Connection Information

  1. Enter Meeting ID: 895800575
  2. Press #

Agenda

  1. GEM Mapping and Simple Tracking Algorithm
  2. Commissioning plan updates from detector systems
  3. Status update from detector owners, calibrations, efficiencies, etc.
  4. Review systems owners and update information
  5. Address system specific points of interest for the Fall 2017 run

Minutes

HMS

12 GeV Drift Chambers

  • Old HMS chambers are out, test stand for 12 GeV chambers are being fit
  • Work to map the HMS dipole is tentatively planned

SHMS

Optics

  • Holly reported that the mapping is finished for nonlinearities
  • HB training is being conducted at 3200 A (80% of 11 GeV maximum)
  • Needs to be retrained at higher current, as noted by Mark and Dave
  • More work is planned after lunch

GEM

  • Latif showed some slides
  • He is currently working on the mapping for the GEM
  • The mapping appears to be working for the most part although more statistics are required in order to populate the x,y position distributions
  • Development for determining the x,y position is ongoing
  • The mapping corrections are implemented and functional
  • Algorithm to handled more complicated situations is required
  • Correlations between the GEM and DCs has been observed
  • It was suggested that the first time bin be utilized for selecting sampling the pedestal
  • Focus should be directed on extracting the pulse time from the data so that it can be correlated to the trigger time
  • Different function for fitting the charge distribution was suggested by Vlad as opposed to the Landau distribution
  • All the data was taken with the SHMS multi crate readout

Drift Chambers

  • Problematic connector is still has problems
  • DC3 won’t be used this spring, old chambers will be used for fall and spring
  • Howard wants to be pinged if chamber is opened again

Hodoscopes

  • Simona reported that 11 bars are done, need to be light leak checked
  • All should be in the test stand on Saturday
    • Then 4 bars will be made
  • Eric reported that the time-walk calibration is 95% complete, although are some discrepancies in the consistency of the fit parameters that are not understood at the moment
  • The corrections will need to go into the decode method of the scintillator classes sooner than later

Aerogel Cerenkov

  • Next week should get spare PMT from CUA, then will need a week to test, install, and check
  • HV module was changed and is behaving nominally